OpenStudy (anonymous):

find the 7th term of a geometric sequence with t1=5 and r=-2

OpenStudy (cwrw238):

nth term of a GS = t1*r^(n-1)

OpenStudy (jdoe0001):

\(\bf \large t_{\color{red}{ n}}=t_1\cdot r^{{\color{red}{ n}}-1} \\ \quad \\ t_1=\textit{1st term in the sequence}\qquad r=\textit{common difference or multiplier/factor} \\ \quad \\ \large t_{\color{red}{ 7}}=t_1\cdot r^{{\color{red}{ 7}}-1}\)